OpenStudy (seratul):

After two years an investment of $800 that is compounded semiannually is worth $920. What is the APY?

OpenStudy (seratul):

@mathmale @satellite73 @zepdrix

satellite73 (satellite73):

you know the formula for this?

OpenStudy (mathnerdz):

^^

OpenStudy (seratul):

Yea. A= Ao(1+r/n)^(n)t

OpenStudy (seratul):

?

OpenStudy (seratul):

@Jamierox4ev3r

pooja195 (pooja195):

@satellite73

OpenStudy (mathmale):

After two years an investment of $800 that is compounded semiannually is worth $920. What is the APY? Initial amount: $800 Final amount: $920 interest rate: unknown number of compounding periods per year: 2 number or years: 2 Insert the knowns into this equation: A= Ao(1+r/n)^(n)t and solve for r.

OpenStudy (seratul):

I got .0711161527 for r. What do I do next?

OpenStudy (sshayer):

\[920=800\left( 1+\frac{ r }{ 2 \times 100 } \right)^{2 \times 2}\] \[\frac{ 920 }{ 800 }=\left( 1+\frac{ r }{ 200 } \right)^4\] \[1+\frac{ r }{ 200 }=\left( \frac{ 23 }{ 20 } \right)^{\frac{ 1 }{ 4 }}=\left( 1.15 \right)^{\frac{ 1 }{ 4 }}\] r=?